OpenStudy (anonymous):

Diane flipped a coin 600 times and it landed heads up 240 times. What is the relative frequency of the coin landing heads up based on the results of this experiment? 4% 6% 40% 60%

OpenStudy (anonymous):

\[240\div600\]

OpenStudy (anonymous):

then multiply that by 100

OpenStudy (anonymous):

percent of x= x/100 so try making denominator as 100 this can be done by dividing it by 6 so divide 240 by 6 240/600 x 6/6 40 /100 so 40 is % or simply 240/600 = x/100 x=40

OpenStudy (anonymous):

am i right

OpenStudy (anonymous):

yes @Dubba2000

OpenStudy (welshfella):

@Dubba2000 - please dont give direct answers

OpenStudy (anonymous):

yea @Dubba2000

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Thank you guys so much!